For the following reaction, if i start this reaction with 153.0 grams of hydrobromic acid and an excess of potassium bi sulfate, a) how many grams of sulfur dioxide gas will I produce? b) if 98.75 g are actually achieved, what is the % yield of the reaction? HBr(aq) + KHSO3(aq) --> KBr(aq)+SO2(g)+H2O(I)  
Answer #1

a)

Balanced reaction

HBr (aq) + KHSO3 (aq) \to KBr (aq) + SO2 (g) + H2O (g)

mole ratio of HBr and SO2 = 1:1

now moles of HBr used

= mass of HBr / molar mass of HBr

= 153.0 (g)  \times1 (mole) 80.91(9)

= 1.89

therefore moles of SO2 can be formed = 1.89

mass of SO2 = moles* molar mass

= 1.89 (mol) \times 64 (\small \frac{g}{mol} )

= 121.02 g

b)


Percent yield

= actual yield x 100 theoretical yield

= 98.75 X 100 121.02

= 81.59 %